OpenStudy (anonymous):

The price of cherries at a farm is $1.38 per pound. Which equation can be used to determine c, the total price of n pounds of cherries?

Mehek (mehek14):

equations?

OpenStudy (anonymous):

c = 1.38 + n c = 1.38 – n c = 1.38n

OpenStudy (anonymous):

I think c, 1.38n

Mehek (mehek14):

`per pound` = multiplication

OpenStudy (anonymous):

yes so C, 1.38n

Mehek (mehek14):

yup

OpenStudy (anonymous):

thank you I have more.

Mehek (mehek14):

oke